OpenStudy (anonymous):

IM DOING applications with polynomials::: 1:) Solve: x4 − 5x3 + 11x2 − 25x + 30 = 0

myininaya (myininaya):

So I tried factoring the first little part.... \[x^4-5x^3+?x^2-?x^2+11x^2-25x+30=0\] \[(x^4-5x^3+?x^2)+(-?x^2+11x^2-25x+30)=0\] \[x^2(x^2-5x+?)+([11-?]x^2-25x+30)=0\] We need to play with that ? a little. What can you replace the question mark in the first trinomial to make it factorable? We are looking at \[x^2-5x+?\]

myininaya (myininaya):

Hello?